### v1

# AudioBench: A Universal Benchmark for Audio Large Language Models

Abstract
We introduce AudioBench, a new benchmark designed to evaluate audio large language models (AudioLLMs). AudioBench encompasses 8 distinct tasks and 26 carefully selected or newly curated datasets, focusing on speech understanding, voice interpretation, and audio scene understanding. Despite the rapid advancement of large language models, including multimodal versions, a significant gap exists in comprehensive benchmarks for thoroughly evaluating their capabilities. AudioBench addresses this gap by providing relevant datasets and evaluation metrics. In our study, we evaluated the capabilities of four models across various aspects and found that no single model excels consistently across all tasks. We outline the research outlook for AudioLLMs and anticipate that our open-source code, data, and leaderboard will offer a robust testbed for future model developments. [^1] https://github.com/AudioLLMs/AudioBench
AudioBench: A Universal Benchmark for Audio Large Language Models
[…]
1 Introduction
Foundation models, built upon large language models (LLMs), have demonstrated strong capability in handling diverse tasks and diverse modalities Bommasani et al. (2021); Achiam et al. (2023); Team et al. (2023). There have been a series of benchmarks proposed recently to provide a holistic evaluation of LLMs Hendrycks et al. (2021); Cobbe et al. (2021a); Wang et al. (2024); Rein et al. (2023), image-enhanced multimodal LLMs Marino et al. (2019); Yue et al. (2023); Padlewski et al. (2024), and even video-enhanced multimodal LLMs Xiao et al. (2021); Li et al. (2023). However, the audio large language models (AudioLLMs) lag, and the performance of their audio interpretation capability is often intractable and not systematically compared in different tasks. The existing evaluation regime does not cover the breadth of their possible use cases.
Although instruction-following audio-language models (or speech-text models) have attracted significant interest from both industry and research communities, their evaluation datasets are notably different. For instance, the Qwen-Audio Chu et al. (2023) model was tested on 12 distinct datasets, whereas the SALMONN Tang et al. (2024) model used 15 different datasets, with only two datasets common between them. Although WavLLM Hu et al. (2024a) has been compared with earlier models, the scope of evaluation tasks remains limited. Another limitation arises from the evaluation set; similar to text-based LLMs Chiang et al. (2024), evaluations predominantly rely on existing datasets and metrics Hendrycks et al. (2021). Given that these models follow instructions, traditional evaluation metrics can not fully meet the requirements and could introduce bias in evaluation. Consequently, it is crucial to develop innovative benchmarks for AudioLLMs. Since their ultimate goal is to handle diverse audio inputs and respond accurately to user queries, the evaluation set should include not only traditional speech tasks but also new datasets and metrics that better reflect real-world applications and assess the models’ ability to follow instructions.
Figure 1: Overview of the datasets in AudioBench.
In this study, we introduce AudioBench as a comprehensive evaluation benchmark specifically designed for general instruction-following audio-language models, which we refer to as AudioLLMs. These models are designed to take comprehension of general audio inputs. These models are designed to comprehend general audio inputs. The tasks should include, but are not limited to, understanding speech semantics, interpreting environmental sounds, and recognizing paralinguistic features, as well as combinations of these elements. Therefore, as shown in Figure 1, we incorporate 8 tasks and 26 datasets covering these aspects, with 7 of the datasets being newly adapted or collected to address gaps in the existing datasets.
For a fair comparison, our benchmark incorporates a variety of prompt templates to evaluate model compatibility with different instructions. Additionally, we vary the input lengths from seconds to minutes to better assess model performance on longer audio sequences. Our test suite includes 400+ hours of audio, distributed across 100k+ samples. The AudioBench toolkit will be made available, enabling the evaluation of future models simply by implementing a customized model inference program.
We also explore evaluation metrics for the open-ended generation of AudioLLMs, which primarily operate as prompt-based, instruction-following models that integrate user queries into their output generation, conditioned on audio inputs. Unlike text-based large language models (LLMs), which can be accessed with multiple-choice questions, AudioLLMs are designed to manage free-format queries. This introduces difficulties in quantitatively assessing their outputs, as using LLMs like GPT-4 as judges can be costly and subject to instability due to version updates. Therefore, we investigate open-source model-as-judge methods and examine their correlation with GPT-4 models. Our findings indicate that LLaMA-3 demonstrates a higher correlation compared to Prometheus-2 Kim et al. (2024) which is specifically fine-tuned for evaluation purposes.
We conducted evaluations on four open-sourced models to offer a comprehensive review of various solutions, including both end-to-end and cascaded models. Our benchmark results indicate that none of the models performed exceptionally well across all criteria, highlighting significant opportunities for future advancements of AudioLLMs with our benchmark. To guide these developments, we also include a detailed discussion on potential future directions for AudioLLMs.
2 Related Work
Multimodal LLM Benchmark. Recent progress in both text-based and multimodal LLMs has spurred the creation of a series of benchmarks designed to monitor and direct their development efficiently. Our research is consistent with earlier benchmarks on multimodal LLMs, primarily concentrating on visual components. There is a lack of comprehensive benchmarks for multimodal LLMs that provide a holistic understanding of audio inputs including speech, ambient sounds, paralinguistic features, etc. When classifying the models targeted for evaluation, we can categorize them into several distinct types: Text-based LLMs, Vision LLMs, and Audio LLMs.
Text-based LLMs present the most fundamental advancements in the past years and serve as the core for developing multimodal systems. Their evaluation is also multifaceted covering aspects such as reasoning Zellers et al. (2019); Cobbe et al. (2021b); Shi et al. (2022), subject knowledge Clark et al. (2018); Hendrycks et al. (2021); Huang et al. (2024b), safety Zhang et al. (2023b); Li et al. (2024b), multilingual capability Wang et al. (2024); Ahuja et al. (2023), and more Chiang et al. (2024); Zheng et al. (2024); Dubois et al. (2024). Text-based LLMs have driven fundamental advancements in recent years Jiang et al. (2023); Achiam et al. (2023); Mesnard et al. (2024) and have significantly contributed to the evolution of vision LLMs that integrate visual understanding with text queries Team et al. (2023). Additionally, a series of multimodal benchmarks specifically tailored for vision LLMs have also been established.
The evaluation of Vision-LLMs is multifaceted, encompassing aspects such as multidisciplinary understanding Yue et al. (2023); Zhang et al. (2024b); Hu et al. (2024b), perception tests Yu et al. (2023); Liu et al. (2023), mathematical reasoning Li et al. (2024a); Zhang et al. (2024a), and video understanding Li et al. (2023); Ning et al. (2023); Liu et al. (2024); Fu et al. (2024). Notably, video understanding benchmarks assess modalities that include both visual and audio elements. However, these benchmarks cannot be directly applied to audio understanding, as they predominantly rely on visual inputs with audio serving merely as a supplementary feature. Consequently, there is an urgent need to standardize benchmarks for Audio LLMs that specifically measure model performance with audio context. Such benchmarks are essential for advancing research and improving the capabilities of models that focus on audio processing and understanding.
Advancements in AudioLLMs. Although no comprehensive evaluation benchmarks currently exist, we observe various multimodal foundation models that enhance speech and audio understanding, utilizing either cascaded methods or integrated multitask optimization. Some models specialize in Automatic Speech Recognition (ASR) Shu et al. (2023); Ma et al. (2024) or other non-speech audio tasks Kong et al. (2024). However, ideal speech-text foundation models often employ multitasking approaches. Notable examples include AudioGPT Huang et al. (2023), SpeechGPT Zhang et al. (2023a), SALMONN Tang et al. (2024), Qwen-Audio Chu et al. (2023), AudioPALM Rubenstein et al. (2023), LTU-AS Gong et al. (2023), VioLA Wang et al. (2023), LLaSM Shu et al. (2023), Pengi Deshmukh et al. (2023), WavLLM Hu et al. (2024a), UniverSLU Arora et al. (2024), and SpeechVerse Das et al. (2024), each focusing on a subset of tasks tailored to their specific use cases. In this study, we focus on three models—WavLLM, Qwen-Audio, and SALMONN—that represent a broader range of task-handling capabilities.
3 AudioBench
3.1 Main Concept
The design of AudioLLMs is to interpret audio content and respond flexibly to user queries, handling tasks from basic speech recognition to complex reasoning. To evaluate these models effectively, we should assess their ability to understand not only speech but also environmental sounds and paralinguistic features such as emotion, gender, and accent. Therefore, our evaluation involves designing benchmarks that cover three key aspects: 1) speech understanding, focusing on the semantic content of audio and including tasks such as Automatic Speech Recognition (ASR), Speech Question Answering (SQA), and Speech Instruction (SI); 2) audio scene understanding, focusing on non-human audio sounds and including tasks such as Audio Question Answering (AQA) and Audio Captioning (AC); and 3) voice understanding, focusing on non-speech but human-related information and including paralinguistic tasks such as Emotion Recognition (ER), Accent Recognition (AR), and Gender Recognition (GR). For each task, we compile existing datasets or develop new ones tailored to the requirements for comprehensive understanding and reasoning. Figure 1 provides an overview of the tasks and the datasets used in the initial setup.
In the current benchmark, our focus is on English audio understanding, where most models have not yet demonstrated satisfactory results. Meanwhile, multilingual understanding, which includes speech translation and code-switching, represents a useful application that could be integrated into the benchmarking datasets as part of an ongoing development effort.
Category Dataset Name # of Samples Total Hours Avg Len (s) Min Len (s) Max Len (s) Audio Understanding ASR LibriSpeech-Clean 2,617 5.40 7.43 1.28 34.96 LibriSpeech-Other 2,935 5.34 6.55 1.47 34.51 CommonVoice-15 16,348 26.95 5.93 1.34 105.67 PeoplesSpeech 32,603 59.20 6.54 1.00 99.91 GigaSpeech 18,650 35.09 6.77 1.00 22.03 Tedlium3 1,142 2.61 8.24 1.07 32.55 Tedlium3-Longform 11 2.61 856.51 340.96 1,596.73 Earning-21 44 39.26 3,212.41 1,097.10 5,740.61 Earning-22 125 119.88 3,452.71 874.73 7,407.05 SQA CN-College-Listen 2,271 13.3 21.09 5.76 137.82 SLUE-P2-SQA5 408 4.5 39.85 13.00 40.0 DREAM-TTS 1,913 18.1 34.14 3.14 261.85 Public-SG-SpeechQA 688 7.6 39.86 15.78 95.71 SI OpenHermes-Audio 100 0.16 5.95 2.04 15.77 ALPACA-Audio 100 0.12 4.32 1.80 8.85 Audio Scene Understanding AQA Clotho-AQA 2,251 14.1 22.59 15.03 29.97 WavCaps-QA 304 0.87 10.28 1.0 30.62 AudioCaps-QA 313 0.86 9.86 3.27 10.00 AC WavCaps 1,730 4.9 10.22 1.00 30.97 AudioCaps 4,400 12.1 9.86 1.74 10.0 Voice Understanding ER IEMOCAP-Emotion 1,002 1.3 4.51 0.75 24.12 MELD-Sentiment 2,610 2.4 3.35 0.13 304.96 MELD-Emotion 2,610 2.4 3.35 0.13 304.96 AR VoxCeleb1-Accent 4,874 11.2 8.27 3.96 69.04 GR VoxCeleb1-Gender 4,874 11.2 8.27 3.96 69.04 IEMOCAP-Gender 1,002 1.26 4.55 0.69 26.77
Table 1: Datasets from the AudioBench are entirely in English and maintain a uniform sampling rate of 16 kHz. Datasets highlighted in gray and green represent newly expanded datasets and those introduced specifically for this study, respectively. Full references can be found in Figure 4.
3.2 Evaluation Setup
From a use case perspective, we expect LLMs to respond to user queries and generate natural responses. Therefore, we evaluate the models in as many free-generation scenarios as possible. For instance, a pair ($Audio$, $Query$) is presented to the model to elicit a ($Model\_Answer$), which is then compared with a reference answer for relevant tasks. This approach contrasts with traditional speech benchmarks like SUPERB Yang et al. (2024) and Dynamic-SUPERB Huang et al. (2024a), where a supervised dataset is specifically fine-tuned for a downstream task and evaluated in a single-task setting.
Therefore, we employed the model-as-judge approach for all tasks except for ASR, where word error rate (WER) is used as the sole metric, and audio captioning, where the METEOR score is utilized as an additional measure. In subsequent experiments, we explored the impact and best practices of the model-as-judge method and validated a reasonable solution.
Moreover, we found that models can be less robust to diverse instruction. This phenomenon also applies to text-based LLMs Zhu et al. (2023); Wang et al. (2024) and we found it even more severe for AudioLLMs as additional modality presents more challenges in the fusion process. Therefore, we propose to leverage multiple instructions to evaluate a single dataset if the query is not diverse enough. This applies to tasks like ASR, ER, AR, GR, and AC. We present this analysis in Section 4.
Figure 2: Sensitivity tests were conducted on three different prompts, revealing that SALMONN is the least robust when faced with varying prompt templates for the same task.
3.3 Tasks
In this section, we discuss the functionality and applicability of each task within this benchmark. For newly proposed datasets, we provide a detailed description of their development process. The detailed statistics are provided in Table 1.
3.3.1 Speech Understanding
Automatic Speech Recognition (ASR). As a fundamental task in speech processing, ASR aims to convert spoken content into text format. It measures the accuracy of speech-to-text conversion and requires robust algorithms capable of understanding and processing a wide range of linguistic nuances and dialects. ASR systems must be equipped to handle diverse environments, speech rates, accents, and background noises, ensuring that they deliver high-quality transcription under various conditions. To test these capabilities, we included 9 datasets, 3 of which contain audio in a long-form format. Current AudioLLMs struggle with long audio files, which may exceed 10 minutes in duration, presenting a potential area for enhancement in future models. For our current evaluations, we segment the long audio into smaller chunks and then reassemble them for assessment if the model naively does not support the original length. Examples can be found at Table 4.
Figure 3: Correlation results for three benchmark datasets across three variations of the model-as-judge approach were analyzed using Spearman’s rank correlation. The findings show that Llama-3-70B-Instruct has the highest correlation with GPT-4.
Speech Question Answering (SQA). The task involves responding to questions based on speech-related audio content. However, there is a lack of suitable datasets. In this part, we involve both monologue understanding and dialogue understanding tasks and curated 3 new datasets. Our collection includes both synthesized and real speech Examples can be found in Table 5.
First, we collected questions from the English listening comprehension section of China’s national college entrance examinations, aimed at assessing students’ ability to understand English in both academic and everyday situations. We manually compiled 271 questions and added them to 2000 questions sourced from Hu et al. (2024a), which includes questions from approximately 130 different exam sets. In line with the free-text QA format, we presented the questions without multiple-choice options. The dataset is named as CN-College-Listen.
DREAM-TTS dataset is built upon text-based dialogue comprehension datasets DREAM Sun et al. (2019). We leverage one SOTA TTS engine - XTTS__\__v2 Casanova et al. (2024) to convert the text input into spoken format. The genders are keep consistent with the 60 diverse speakers randomly selected.
For the Public-SG-SpeechQA dataset, we chose four public speaking videos from Singaporean politicians, accompanied by clean transcriptions. These transcriptions were manually segmented, and five questions per segment were generated using LLMs. Each question and its corresponding reference answer were manually reviewed to ensure their validity, where 30% of the samples are discarded. Ultimately, this process resulted in the collection of 688 speech question-answer pairs for evaluation.
Speech Instruction (SI). This task evaluates whether the model can directly follow instructions provided through speech input, mirroring a natural human-computer interaction. Specifically, the question is delivered via audio, and the model is expected to understand and generate responses in the appropriate text format. This approach offers unique benefits. For instance, by incorporating paralinguistic information such as emotions, the model can adapt its responses based on the user’s emotional state. However, finding suitable testing examples has been a challenge. In this study, we use a TTS system Casanova et al. (2024) to convert an existing instruction-following dataset into the spoken format. We then involve humans in selecting instances where the speech is accurately synthesized and the content is suitable as spoken instructions. As a result, 10.5% samples are kept and the examples can be found in Table 8.
Dataset Name Metrics AudioLLMs Whisper+Llama3 SALMONN Qwen-Audio WavLLM Audio Understanding LibriSpeech-Clean WER 45.28 2.20 2.05 1.83 LibriSpeech-Other 33.76 4.40 4.88 3.71 CommonVoice-15 44.44 11.30 14.84 9.89 PeoplesSpeech 35.05 30.59 37.51 14.55 GigaSpeech 14.42 13.18 16.02 9.52 Tedlium3 13.21 4.28 5.43 3.81 Tedlium3-Longform 25.88 34.16 46.51 4.94 Earning-21 31.81 39.39 65.64 11.77 Earning-22 35.87 54.37 67.22 15.60 CN-College-Listen M.J.70B 1.95 2.15 2.51 2.96 SLUE-P2-SQA5 3.95 3.84 4.16 4.19 DREAM-TTS 1.79 1.91 2.50 2.82 Public-SG-SpeechQA 2.77 2.97 2.98 3.25 OpenHermes-Audio 1.05 1.14 1.07 2.82 ALPACA-Audio 0.85 1.12 1.15 3.37 Audio Scene Understanding Clotho-AQA M.J.70B 2.75 2.88 2.16 1.40 WavCaps-QA 2.31 1.97 1.34 0.83 AudioCaps-QA 2.39 2.41 1.34 0.77 WavCaps 1.25 1.41 0.49 0.23 AudioCaps 1.83 2.11 0.50 0.24 WavCaps METEOR 21.36 21.75 9.78 14.52 AudioCaps 22.79 22.91 6.70 7.15 Voice Understanding IEMOCAP-Emotion M.J.70B 2.33 1.60 2.43 1.85 MELD-Sentiment 1.82 1.72 1.65 2.26 MELD-Emotion 1.60 1.46 1.45 1.88 VoxCeleb1-Accent 1.50 2.25 1.87 1.19 VoxCeleb1-Gender 3.98 3.59 2.96 2.87 IEMOCAP-Gender 2.02 2.55 2.01 2.51
Table 2: Main evaluation results of various AudioLLMs and one cascade model. M.J.70B refers to Model-as-Judge with Llama-3-70B-Instruct model (Score is from 0-lowest to 5-highest). METEOR scores are common metrics for audio captioning Banerjee and Lavie (2005). WER is the word error rate for ASR tasks. When outperforming end-to-end models, the Whisper+Llama3 model is marked with underline.
3.3.2 Audio Scene Understanding
Audio Question Answering (AQA). This task focus more on the understanding of environmental contexts. To assess the capability to follow instructions in these scenarios, we refined a test set from Clotho-AQA, retaining only the samples with high confidence levels. Simultaneously, we developed the WavCaps-QA and AudioCaps-QA datasets, each featuring over 300 diverse questions accompanied by human-verified questions and answers. Examples can be found in Table 6.
Audio Captioning (AC). This task involves generating captions (descriptions) for an audio clip. To assess this task, we incorporate two popular datasets: WavCaps and AudioCaps. The examples can be found at Table 9.
3.3.3 Voice Understanding
Emotion Recognition (ER). Emotion is a critical paralinguistic element that can be conveyed through human speech or non-speech content El Ayadi et al. (2011). It plays a vital role in making communications interpretable for listeners. Three datasets are incorporated and the examples can be found at Table 7.
Accent Recognition (AR). Accent, a significant paralinguistic feature, has been overlooked in previous evaluations. We have included metadata from the VoxCeleb1 dataset and tasked the model with predicting the speaker’s likely origin based on their accent. Examples can be found at Table 10.
Gender Recognition (GR). Gender could be recognized based on vocal characteristics. This task tests the model’s sensitivity to gender-specific acoustic patterns. We have prepared two datasets and the examples can be found at Table 11.
4 Results and Analysis
Models. For comparative analysis, we selected three representative AudioLLMs: SALMONN Tang et al. (2024), Qwen-Audio Chu et al. (2023), and WavLLM Hu et al. (2024a). These models are among the most capable multi-task speech-text models available, designed to simultaneously manage a variety of tasks. Additionally, we include a cascade model named Whisper+Llama3, which processes data in a pipeline manner. First, transcriptions are extracted using the Whisper-larg-v3 Huang and Tsai (2023) model, and then these transcriptions along with user queries are input into the Llama-3-8B-Instuct model to generate responses. While this model lacks the ability to comprehend rich audio content, relying solely on transcriptions for context, it demonstrates strong performance in certain speech-intensive tasks and serves as a robust baseline.
Robustness to Diverse Queries. In evaluation, we found that AudioLLMs could behave differently to different prompt templates. Even though this also applies to text-based LLMs, it is even more severe in AudioLLMs as the two modalities are not aligned in the first stage and aggressive alignment would hurt the generalizability to different prompts. Hu et al. (2024a) also spot such issues but only limited experimental studies are presented. To study such effect, we did a more thorough evaluation in this case with 2 models (SALMONN and Qwen-Audio) on 3 ASR datasets (LirbriSpeech-Clean, CommonVoice, Tedium3).
Figure 2 illustrates the results. We can witness that SALMONN can effectively perform ASR tasks and achieve performance comparable to the Qwen-Audio model on the LibriSpeech-Clean and Tedium3 tasks, provided the right prompts are used. Specifically, prompt 3, "Decode the audio and give me the written transcripts," triggers the SALMONN model to conduct phoneme recognition on the LibriSpeech and Tedium3 datasets. For the CommonVoice dataset, SALMONN tends to perform speech translation on a significant number of samples, which adversely affects its performance on the Word Error Rate (WER) metric. This suggests that SALMONN may be overly tuned to speech features (tokens) and not sufficiently responsive to the prompts. In contrast, Qwen-Audio’s performance remains stable across all three prompt templates. Consequently, to minimize variability and thoroughly test model robustness, we incorporate at least 20 diverse prompt templates into our evaluation framework for the tasks without diverse prompts originally.
Model-as-Judge Comparison. In many datasets, we use an open-ended generation style, which is considered the most natural method of interaction. However, this style introduces challenges for evaluation, as traditional metrics are not easily adaptable. To address this, we employ a model-as-judge approach to assess model responses against reference answers and questions. Yet, there are ongoing challenges in selecting appropriate models to ensure both accessibility and robustness.
The most frequently used judge model is GPT-4. However, GPT-4 is closed-source, subject to ongoing updates, and often incurs costs that may not be feasible for the wider community. To explore alternatives, we examined 3 open-source models: Llama-3-8B-Instruct, Llama-3-70B-Instruct, and Prometheus-2 Kim et al. (2024). Prometheus-2 is specifically fine-tuned for response grading, while Llama-3-Instruct is a general-purpose open-source model.
To identify the most suitable open-source model for evaluation tasks, we use all samples from the following three datasets: CN-College-Listen, Clotho-AQA, and VoxCeleb1-Accent, to evaluate their correlation with GPT-4. The three datasets cover speech understanding, audio scene understanding, and voice understanding. Initially, we get the model outputs from the SALMONN model. These model outputs, along with golden answers and questions, are then fed into the four judgment models to obtain judgment scores. We subsequently calculated the Spearman’s rank correlation to compare the effectiveness of the different judgment models. The evaluation template can be found in Table 3. The results are detailed in Figure 3.
The Llama-3-70B-Instruct model exhibits the highest correlation with GPT-4-as-a-judge (gp4-turbo-0409). The correlation scores exceeding 0.85 across all three datasets indicate a "very strong correlation" with GPT-4. We attribute this to Llama-3-70B-Instruct’s robust generalizability. In contrast, Prometheus-2, though fine-tuned from the Mistral model and further adapted for scoring purposes, does not effectively compensate for the limitations of its base model. Therefore, in subsequent experiments, we will use Llama-3-70B-Instruct as the judging model (score 0 to 5), as it provides transparency and offers ease of adaptability. It is important to note that using a model-as-judge is not an ideal solution for grading open-ended generations and has several shortcomings. The issue of accurately grading free-form generations remains an unresolved challenge.
Main Results Table 2 detailed the overall results across 26 datasets for 4 models, revealing that no single model consistently outperforms the others on all tasks. The SALMONN model is sensitive to varying instructions for ASR tasks, while Qwen-Audio and WavLLM demonstrate robust ASR capabilities. However, these AudioLLMs struggle with long-form ASR tasks without adequate context. Additionally, the training and pre-training datasets for AudioLLMs are significantly smaller than those used for other acoustic models like Whisper, which limits their generalizability to unnatural truncation.
For speech-intensive tasks such as SQA and SI, the cascade model Whisper+Llama3 exhibits superior performance. This effectiveness stems from the Whisper model’s robust recognition capabilities and Llama’s strong reasoning abilities, which together efficiently handle queries primarily contained in the verbal content. In contrast, the modality fusion process in AudioLLMs may distort speech content, highlighting an area for potential improvement.
For tasks involving paralinguistic features and non-verbal sounds, AudioLLMs generally outperform cascade models, although they do not always yield satisfactory results. An exception is found in sentiment and emotion recognition tasks, where some emotions can be directly inferred from the speech semantics. Overall, cascade models struggle with understanding non-verbal content, highlighting the need for the more robust development of AudioLLMs to handle these complexities better.
In comparing three AudioLLMs, we observed that models perform better when exposed to related tasks during training. A significant example is WavLLM, which specializes in speech content integration. It excels in SQA tasks but lacks exposure to non-spoken scenarios such as audio captioning. Consequently, it underperforms in audio scene understanding and vocal understanding tasks, demonstrating that its instruction-following capabilities are not readily generalizable without specifically aligned training samples.
5 Research Outlook
Long Audio Processing and Understanding: The current benchmarks primarily assess the understanding capabilities of AudioLLMs using audio clips of limited duration (within minutes). However, extending the capability to process longer audio can open up broader applications such as meeting summarization and sequential event understanding. In text-based LLMs, long-sequence processing has advanced rapidly. By embedding speech content as tokens, this capability could be effectively explored.
Multi-round Query Handling: The ability to manage multi-round queries is still limited in open-source models. Enhancing this feature would allow for more dynamic interactions where each query could involve different modalities such as images or audio, making the models more versatile in practical applications.
Multilingual Capabilities, Code-Switching, and Dialects: Expanding the linguistic capabilities of AudioLLMs to handle multiple languages, code-switching, and various dialects is crucial. This enhancement would improve the models’ applicability across diverse linguistic and cultural contexts, making them more effective for global use.
Speech Generation: Developing more sophisticated speech generation capabilities within AudioLLMs would enable more natural and engaging human-computer interactions. This includes improving the models’ ability to generate speech that not only conveys information but also mimics human-like intonations and rhythms.
6 Conclusion
In this work, we introduce the first comprehensive evaluation benchmark for AudioLLMs, named AudioBench, which includes 8 tasks and 26 datasets. Our report includes a detailed analysis of data selection, metric design, and the evaluation pipeline. We emphasize the critical factors for benchmark design and model development. We anticipate that our proposal will stimulate further progress in multimodal foundation models.
Limitations
First, the current AudioBench exclusively includes English datasets. However, multilingual capabilities and code-switching are crucial for comprehensive speech understanding and generation. We plan to expand the benchmark to incorporate these aspects in future iterations.
Second, similar to text-based LLMs, evaluating free-style generation is challenging and demands robust metrics or models to serve as judges. Traditional metrics fall short in zero-shot scenarios, and assessing the correctness of output responses is complex. Thus, the development of suitable evaluation metrics, particularly for audio inputs, is crucial.
Third, as end-to-end models, AudioLLMs typically involve large model sizes, which result in longer inference times. In this benchmark, our focus has been primarily on accuracy rather than efficiency. Moving forward, it will be important to consider inference speed and the deployment environment to provide a comprehensive evaluation of these models during the deployment process.
Acknowledgement
We are grateful to Ziyi Xu (NUS, Singapore) and Anh Thuc Nguyen (UNC, USA) for their kind contributions to annotations and quality verification.
Appendix A Appendix
In Figure 4, we present the structure of the tasks and datasets included in our benchmark. In this section, we aim to provide a detailed description of the newly curated datasets along with their construction processes.
CN-College-Listen. This is a zero-shot testing set compiled from English comprehension questions extracted from college exam papers. Initially, we downloaded audio files and corresponding exam papers from public websites. These audio files were segmented according to the questions in the exam papers. Typically, an audio segment may relate to multiple questions depending on the sections of the exam paper. Each segmented audio is paired with its corresponding questions to formulate the test items. Originally, the questions in the exam papers are formatted as multiple-choice; however, for the application with AudioLLMs, it is more natural to frame them as open-ended questions. We retain the correct choice as the reference answer. In total, we collected 271 audio-question-answer triples and supplemented these with another 2000 questions from Hu et al. (2024a) to create the final dataset. This dataset requires a precise capture of audio content, including both monologues and dialogues.
DREAM-TTS. The evaluation set for spoken dialogue understanding remains limited. DREAM is a dataset featuring human-annotated questions and answers related to a dialogue. We used text-to-speech (TTS) technology to convert the original text dialogues into spoken format, adhering to gender information to closely mimic real-life scenarios. The original questions and answers were retained to create dialogue-question-answer triples.
Public-SG-SpeechQA. To evaluate real speech content, we selected four videos from public speeches in Singapore. These speeches vary in length from 20 minutes to an hour, and clean transcripts are available online. We manually segmented the speeches based on topic transitions. Using the Llama-2-7B-Chat model, we generated question-and-answer pairs for each segment. These pairs underwent a subsequent round of human review to ensure they were concrete and relevant. Ultimately, we compiled 688 speech-question-answer triples, with each speech segment corresponding to multiple questions and answers.
OpenHermes-Audio and ALPACA-Audio. Since effective speech instruction datasets are lacking, we utilized an open-source instruction tuning set designed for text-based LLMs and converted these instructions into spoken format using text-to-speech (TTS) technology. Human reviewers then verified these converted samples, filtering out approximately 90% and retaining only 10% for our test set. This filtering process ensures the readability, naturalness of the spoken instructions, and correctness of the reference answers. Figure 5 and 6 show the annotation platform we developed using Streamlit for data annotation for these two datasets.
WavCaps QA and AudioCaps QA. Figure 7 illustrates the data annotation platform we developed for the WavCaps and AudioCaps datasets. Initially, we used the Llama-3-8B-Instruction model to generate questions and answers from the provided captions. Each generated sample then went through another round of human annotation to refine the questions and ensure the validity of the test set. Answers were also revised as needed. Through this process, we compiled over 300 questions for each dataset to assess the models’ performance effectively.
Model-as-Judge Templates. Table 3 the scoring templates used for the model-as-judge approach. The scoring range is from 0 to 5, except for Prometheus2, where the range is from 1 to 5.
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw, rounded corners,align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=1pt, , where level=1text width=5em,fill=blue!10, where level=2text width=5em,font=,fill=pink!30, where level=3font=,yshift=0.26pt,fill=yellow!20, [AudioBench
Benchmark for AudioLLMs,fill=green!20 [Automatic Speech Recognition (ASR),text width=13em [LibriSpeech-Clean Panayotov et al. (2015),text width=14.5em] [LibriSpeech-Other Panayotov et al. (2015),text width=14.5em] [CommonVoice Ardila et al. (2020),text width=13.5em] [PeoplesSpeech Galvez et al. (2021),text width=13.5em] [Earning-21 Del Rio et al. (2021),text width=13.5em] [Earning-22 Del Rio et al. (2022),text width=13.5em] [GigaSpeech Chen et al. (2021),text width=13.5em] [Tedium3 Rousseau et al. (2012),text width=13.5em] [Tedium3-Longform Rousseau et al. (2012),text width=15em] ] [Speech Question Answering (SQA),text width=12em [CN-College-Listen Hu et al. (2024a),text width=18em] [SLUE-P2-SQA5 Shon et al. (2022),text width=13.5em] [DREAM-TTS Sun et al. (2019),text width=13.5em] [Public-SG-SpeechQA,text width=13.5em] ] [Audio Question Answering (AQA),text width=12em [Clotho-AQA Lipping et al. (2022),text width=13.5em] [WavCaps-QA Mei et al. (2023),text width=13.5em] [AudioCaps-QA Kim et al. (2019),text width=13.5em] ] [Emotion Recognition (ER),text width=9em [IEMOCAP-Emotion Busso et al. (2008),text width=13.5em] [MELD-Sentiment Poria et al. (2019),text width=13.5em] [MELD-Emotion Poria et al. (2019),text width=13.5em] ] [Speech Instruction (SI),text width=8em [OpenHermes-Audio Teknium (2023),text width=15em] [ALPACA-Audio Taori et al. (2023),text width=15em] ] [Audio Captioning (AC),text width=8em [WavCaps Mei et al. (2023),text width=13.5em] [AudioCaps Kim et al. (2019),text width=13.5em] ] [Accent Recognition (AR),text width=8.5em [VoxCeleb-Accent Nagrani et al. (2017),text width=13.5em] ] [Gender Recognition (GR),text width=8.5em [VoxCeleb-Gender Nagrani et al. (2017),text width=13.5em] [IEMOCAP-Gender Busso et al. (2008),text width=13.5em] ] [Future Tasks,text width=7em [Audio Generation Tasks,text width=9em [Text-to-Speech (TTS)] [Voice Conversion (VC)] [Speech Segmentation (SS)] ] [Multilingual Understanding,text width=10em [Speech Translation (ST) & others.]] [Multiple Audio Understanding,text width=10.5em [Speaker Verification (SV) & others.] ] [etc.,text width=5em] ] ]
Figure 4: Structure of AudioBench datasets.
Dataset Context Instruction (Sample) Answer LibriSpeech-Clean [Uncaptioned image] NO I WASN’T THINKING OF THAT I need the spoken content written down, please. NO I WASN’T THINKING OF THAT LibriSpeech-Other [Uncaptioned image] THE HISTORY OF THE HOUSE IS PLAIN NOW I’m asking for a transcription of the spoken words into text. THE HISTORY OF THE HOUSE IS PLAIN NOW CommonVoice [Uncaptioned image] This is Jon Davis. Could you write out what was said? This is Jon Davis. PeoplesSpeech [Uncaptioned image] that’s where you have a lot of windows in the south no actually that’s passive solar and passive solar is something that was… Could you transcribe what you hear? that’s where you have a lot of windows in the south no actually that’s passive solar and passive solar is something that was… GigaSpeech [Uncaptioned image] I DO THINK THAT FAST FASHION HAS KIND OF ALLOWED EVERYONE TO HAVE ACCESS TO AESTHETICALLY THOUGHTFUL CLOTHING. Capture the speech in written format, please. I DO THINK THAT FAST FASHION HAS KIND OF ALLOWED EVERYONE TO HAVE ACCESS TO AESTHETICALLY THOUGHTFUL CLOTHING <PERIOD> Earning-21 [Uncaptioned image] TOKEN GOOD MORNING EVERYONE AND WELCOME TO THE NEXTERA ENERGY INC AND NEXTERA ENERGY PARTNERS… Please document the spoken language. TOKEN GOOD MORNING EVERYONE AND WELCOME TO THE NEXTERA ENERGY INC AND NEXTERA ENERGY PARTNERS… Earning-22 [Uncaptioned image] TOKEN GOOD DAY LADIES AND GENTLEMEN AND WELCOME TO THE CD PROJEKT GROUP FINANCIAL RESULTS FOR H1 2021 CONFERENCE CALL TODAY’S CALL … Document the conversation by writing it down. TOKEN GOOD DAY LADIES AND GENTLEMEN AND WELCOME TO THE CD PROJEKT GROUP FINANCIAL RESULTS FOR H1 2021 CONFERENCE CALL TODAY’S CALL … Tedium3 [Uncaptioned image] one day los angeles times columnist steve lopez was walking along the streets … Please transcribe the speech. one day los angeles times columnist steve lopez was walking along the streets … Tedium3-Longform [Uncaptioned image] i ’d like to share with you a discovery that i made a few months ago while writing an article for italian wired i always keep my thesaurus handy whenever … Please transcribe the speech. i ’d like to share with you a discovery that i made a few months ago while writing an article for italian wired i always keep my thesaurus handy whenever …
Table 4: Examples from Automatic Speech Recognition (ASR) datasets
Dataset Context Instruction Answer CN-College-Exam-Listening [Uncaptioned image] F: Excuse me, this is the address, where do I find it? M: All right, you need a street map, here is one, and I will show you where it is. What does the woman want to do? Find a place. SLUE-P2-SQA5 [Uncaptioned image] 1 Climate Stephens City is located in the humid subtropical climate zone ( Köppen climate classification : Cfa ) … Which regions have temperate climates? mid-atlantic DREAM-TTS [Uncaptioned image] F: The movie next Tuesday has been canceled due to lack of interest. M: what do you mean? F: Well by last night only a few tickets been sold. What can we conclude about the movie? The movie will not be shown. Public-SG-SpeechQA [Uncaptioned image] Today, speaking to a roomful of economists, I am inclined to confine myself to talk about markets and price, dollars … How can economics help solve complex healthcare challenges, as mentioned by the speaker? Economics can help solve complex healthcare …
Table 5: Examples from Speech Question Answering (SQA) datasets
Dataset Context Instruction Answer Clotho-AQA [Uncaptioned image] (Wave sound) Are there waves? yes WavCaps-QA [Uncaptioned image] (Electronic Music playing) What type of sound is being played? The sound being played is music. AudioCaps-QA [Uncaptioned image] (Mechanical vibration sound) What type of object or equipment is likely to produce a constant rattling noise and sharp vibrations? A loose or worn-out bolt or screw on a machine or equipment is likely to produce a constant rattling noise and sharp vibrations.
Table 6: Examples from Audio Question Answering (AQA) datasets
Dataset Context Instruction (Sample) Answer IEMOCAP-Emotion [Uncaptioned image] Thank you. How would you interpret the speaker’s emotional state from their speech? From the speaker’s speech, it seems they are in a sad state. MELD-Sentiment [Uncaptioned image] Yeah, I’m not in that. How would you interpret the speaker’s sentiment state from their speech? From the speaker’s speech, it seems they are in a neutral sentiment state. MELD-Emotion [Uncaptioned image] Yeah, I’m not in that. What do you think the speaker is feeling based on their speech patterns (e.g., pace, pitch, volume)? Based on the speaker’s speech patterns, it seems like they are feeling neutral.
Table 7: Examples from Emotion Recognition (ER) datasets
Dataset Context Instruction (Sample) Answer OpenHermes-Audio [Uncaptioned image] Pretend to be Ghost, expressing frustration to Soap that they’re no closer to finding the elusive enemy leader after weeks of searching. Please follow the instruction in the speech It feels like we’ve been chasing ghosts, Soap. After all these weeks, we’re still no closer to finding the bastard. It’s bloody infuriating. ALPACA-Audio [Uncaptioned image] identify the type of the sentence: she can play guitar. Please follow the instruction in the speech The type of sentence is a declarative sentence.
Table 8: Examples from Speech Instruction (SI) datasets
Dataset Context Instruction (Sample) Answer WavCaps [Uncaptioned image] (Electronic Music playing) Detail the ambient sounds included in the audio. A sound is playing. AudioCaps [Uncaptioned image] (Mechanical vibration sound) Describe the environmental sounds in the audio. Constant rattling noise and sharp vibrations
Table 9: Examples from Audio Captioning (AC) datasets
Dataset Context Instruction (Sample) Answer VoxCeleb-Accent [Uncaptioned image] because I was trying to get away from her, but where did you go? what did you do? From which country does the speaker hail? From the audio, I guess the speaker is from USA.
Table 10: Examples from Accent Recognition (AR) datasets
Dataset Context Instruction (Sample) Answer VoxCeleb-Gender [Uncaptioned image] because I was trying to get away from her, but where did you go? what did you do? Is there information available on the speaker’s gender? The speaker is a female. IEMOCAP-Gender [Uncaptioned image] For God’s sake, Augie. It’s- Grow up, we’re not going to see the grunion. Based on what you heard, can you identify whether the speaker is male or female? Based on the auditory cues, it sounds like the speaker is a male.
Table 11: Examples from Gender Recognition (GR) datasets
Figure 5: OpenHermes-Audio Labeling Platform.
Figure 6: ALPACA-Audio Labeling Platform.
Figure 7: Labeling Platform for AudioCaps QA and WavCasp QA.
Dataset Name Metrics AudioLLMs Whisper+Llama3 SALMONN Qwen-Audio WavLLM Audio Understanding LibriSpeech-Clean WER 45.28 2.20 2.05 1.83 LibriSpeech-Other 33.76 4.40 4.88 3.71 CommonVoice-15 44.44 11.30 14.84 9.89 PeoplesSpeech 35.05 30.59 37.51 14.55 GigaSpeech 14.42 13.18 16.02 9.52 Tedlium3 13.21 4.28 5.43 3.81 Tedlium3-Longform 25.88 34.16 46.51 4.94 Earning-21 31.81 39.39 65.64 11.77 Earning-22 35.87 54.37 67.22 15.60 CN-College-Listen M.J.8B 1.57 1.63 1.95 2.32 SLUE-P2-SQA5 3.09 3.03 3.08 3.26 DREAM-TTS 1.46 1.35 1.86 2.23 Public-SG-SpeechQA 2.18 2.42 2.18 2.51 OpenHermes-Audio 2.15 2.12 1.72 2.34 ALPACA-Audio 1.72 2.19 1.71 2.39 Audio Scene Understanding Clotho-AQA M.J.8B 2.08 2.26 1.76 1.28 WavCaps-QA 1.90 1.58 1.03 0.95 AudioCaps-QA 1.94 1.93 1.06 1.05 WavCaps 1.18 1.16 0.51 0.65 AudioCaps 1.47 1.39 0.49 0.56 WavCaps METEOR 21.36 21.75 9.78 14.52 AudioCaps 22.79 22.91 6.70 7.15 Voice Understanding IEMOCAP-Emotion M.J.8B 1.59 1.00 1.71 1.44 MELD-Sentiment 1.51 1.21 1.30 1.93 MELD-Emotion 1.16 1.02 1.04 1.47 VoxCeleb1-Accent 1.52 1.59 1.38 0.95 VoxCeleb1-Gender 2.58 2.74 2.41 1.67 IEMOCAP-Gender 1.35 1.95 1.54 1.56
Table 12: Main evaluation results of various AudioLLMs and one cascade model. M.J.8B refers to Model-as-Judge with Llama-3-8B-Instruct model (Score is from 0-lowest to 5-highest). METEOR scores are common metrics for audio captioning Banerjee and Lavie (2005). WER is the word error rate for ASR tasks. When outperforming end-to-end models, the Whisper+Llama3 model is marked with underline.